# Can We Validate Counterfactual Estimations in the Presence of Unknown Network Interference?

### v1

@@ 9.4 Batch-level State Evolution
    For all $i$ and $t\in[T]_{0}$, as $N\rightarrow\infty$, we have
    $\;\mathbb{E}_{p_{y}^{\mathcal{S}}(N)\times p_{x}^{\mathcal{S}}(N)\times p_{w}^{\mathcal{S}}(N)\times p_{\mu^{i}}(N)\times p_{\mu_{t}^{i}}(N)\times p_{\mu}^{\mathcal{S}}(N)\times p_{\mu_{t}}^{\mathcal{S}}(N)}\big{[}{\|Y_{0}{},{\vec{X}},\vec{W},{M}^{i},{M}_{t}^{i},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\|}^{k}\big{]}$
    $\rightarrow$ $\;\mathbb{E}_{p_{y}^{\mathcal{S}}\times p_{x}^{\mathcal{S}}\times p_{w}^{\mathcal{S}}\times p_{\mu^{i}}\times p_{\mu^{i}_{t}}\times p_{\mu}^{\mathcal{S}}\times p_{\mu_{t}}^{\mathcal{S}}}\big{[}{\|Y_{0}{},{\vec{X}},\vec{W},{M}^{i},{M}_{t}^{i},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}}\|}^{k}\big{]}.$
@@
We can replace the second part of Assumption 9.4-(iv) by assuming that $p_{\mu^{i}}$ is the same across all units, and similarly, $p_{\mu^{i}_{t}}$ is identical for all units.
Remark 9.5
We can drop Assumptions 9.4 and 9.4-(v) by confining the functions $g_{t}$ and $h_{t}{}$ to be bounded and continuous.
@@
In Assumption 9.4, when $\mathcal{S}$ represents the entire experimental population, we omit the superscript $\mathcal{S}$ from all notations.
Remark 9.7 (Time-dependent covariates)
For each time $t=0,1,\ldots,T$, let ${\cal X}_{t}\in\mathbb{R}^{n^{\text{x}}_{t}\times N}$ denote the time-dependent covariate matrix, where the $i^{th}$ column corresponds to the covariates for unit $i$ at time $t$. We extend the covariate matrix $\bm{X}$ by incorporating these time-dependent covariates, resulting in a new $(n^{\text{x}}+n^{\text{x}}_{0}+\ldots+n^{\text{x}}_{T})\times N$ covariate matrix in our model. The functions $g_{t}$ and $h_{t}{}$ are then modified accordingly to reference the appropriate portion of this extended covariate matrix at each time period. Along this, we exclude the noise vectors (i.e., $\vec{\epsilon}^{\;}_{t}$) from subsequent discussions on the potential outcome specification Eq. (15).
Assumption 9.4 is a standard assumption in statistical theory, ensuring that the empirical distributions of system attributes remain stable and do not diverge as the sample size increases. This assumption holds, for example, when units’ attributes $\left\{(Y^{i}_{0}{},{\vec{X}}^{i},\vec{W}^{i},\vec{\mu}^{\;i\cdot},\vec{\mu}^{\;i\cdot}_{t})\right\}_{i}$ follow an i.i.d. distribution with finite moments of order $k$. Moreover, a wide range of treatment assignments satisfies the conditions of Assumption 9.4, including cases where the support of $\pi$ is bounded, such as the Bernoulli design. By imposing such conditions, we ensure the reliability of estimation by keeping the experimental design moments finite and manageable.
@@
$\tilde{\nu}_{1}^{\mathcal{S}}{}$ $=\mathbb{E}\left[({M}^{\mathcal{S}}+{M}_{0}^{\mathcal{S}})g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}\right],\quad\tilde{\rho}_{1}{}=(\sigma+\sigma_{0})\mathbb{E}\left[g_{0}\big{(}Y_{0}{},\vec{W},{\vec{X}}\big{)}^{2}\right],\quad H^{\mathcal{S}}_{0}{}=h_{0}{}\big{(}Y^{\mathcal{S}}_{0}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)},$ (27)
@@
  • ${M}\sim p_{{M}}$ and ${M}_{t}\sim p_{{M}_{t}}$ represent the weak limits of interference elements at the population level, as specified in Assumption 9.4. Similarly, the corresponding subpopulation quantities are denoted by ${M}^{\mathcal{S}}\sim p_{{M}}^{\mathcal{S}}$ and ${M}_{t}^{\mathcal{S}}\sim p_{{M}_{t}}^{\mathcal{S}}$;
  • and independent from all of them, $Z_{t}$ follows a standard Gaussian distribution.
The following theorem characterizes the distribution of units’ outcomes $Y^{i}_{1}{},\ldots,Y^{i}_{t+1}{}$ within the large sample asymptotic, based on the BSE equations outlined in Eq. (27).
@@
Fixing $k\geq 2$, consider the sequence of units’ attributes $\left\{(Y^{i}_{0}{},{\vec{X}}^{i},\vec{W}^{i},\vec{\mu}^{\;i\cdot},\vec{\mu}^{\;i\cdot}_{t})\right\}_{i,t}$ and suppose that Assumptions 3.3-9.4 hold. Then, in view of the BSE equations given in Eq. (27), for any function $\psi\in{{\cal C}{\cal P}}(k)$, we have,
@@
=$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{T}^{\mathcal{S}}{}+\tilde{\rho}_{T}{}Z_{T}+H^{\mathcal{S}}_{T-1}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}\Big{]},$
where $Z_{t}\sim{\cal N}(0,1),\;t=1,\ldots,T,$ independent of $\vec{W}^{\mathcal{S}}\sim p_{w}^{\mathcal{S}}$ and ${\vec{X}}^{\mathcal{S}}\sim p_{x}^{\mathcal{S}}$. In addition, for any $t=1,\ldots,T$, the random variables $Z_{t}$ and $H_{t-1}{}$ are independent.
@@
To derive the result in the second part of Theorem 4.2 from Theorem 9.8, we note that when $\mathcal{S}$ depends solely on the treatment allocation, which is independently distributed from all other variables, $\mathcal{S}$ effectively functions as a random sample from the experimental population. As a result, the distributions of $Y^{\mathcal{S}}_{0}{}$, ${\vec{X}}^{\mathcal{S}}$, ${M}^{\mathcal{S}}$, and ${M}_{t}^{\mathcal{S}}$ are equivalent to those of $Y_{0}{}$, ${\vec{X}}$, ${M}$, and ${M}_{t}$, respectively. This equivalence follows from the fact that the empirical distribution of quantities in a random sample converges to the empirical distribution of the entire population.
@@ 9.5 Conditioning Technique
$\bm{Q}_{t}:=\left[\vec{U}_{0}{}\Big{|}\vec{U}_{1}{}\Big{|}\ldots\Big{|}\vec{U}_{t-1}{}\right],\quad\quad\bm{R}_{t}:=\left[\vec{\Upsilon}_{1}{}-{\bf B}_{0}\vec{U}_{0}{}\Big{|}\ldots\Big{|}\vec{\Upsilon}_{t}{}-{\bf B}_{t-1}\vec{U}_{t-1}{}\right].$ (29)
According to Eq. (29), $\bm{Q}_{t}$ and $\bm{R}_{t}$ are matrices with columns of $\vec{U}_{s-1}{}$ and $\vec{\Upsilon}_{s}{}-{\bf B}_{s-1}\vec{U}_{s-1}{}$, when $s=1,\ldots,t$, respectively. Then, we denote by $\vec{U}^{\parallel}_{t}{}$ the projection of $\vec{U}_{t}{}$ onto the space generated by the columns of $\bm{Q}_{t}$ and define $\vec{U}^{\perp}_{t}{}=\vec{U}_{t}{}-\vec{U}^{\parallel}_{t}{}$. We also define $\vec{\gamma}_{t}=(\gamma^{0}_{t},\gamma^{1}_{t},\ldots,\gamma^{t-1}_{t})^{\top}$ such that
@@
Now, note that the available observation at any time $t$ implicitly reveals information about the fixed interference matrix ${\bf A}$. To manage this intricate randomness, we define ${\cal G}_{t}$ as the $\sigma$-algebra generated by $\vec{Y}_{0}{},\vec{Y}_{1}{},\ldots,\vec{Y}_{t}{}$, $\vec{\Upsilon}_{1}{},\ldots,\vec{\Upsilon}_{t}{}$, ${\bf M}_{t}$, ${\bf B}_{0},\ldots,{\bf B}_{t-1}$, $\bm{W}{}$, and $\bm{X}$. We then compute the conditional distribution of ${\bf A}$ given ${\cal G}_{t}$. In this framework, conditioning on ${\cal G}_{t}$ is equivalent to conditioning on the event ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$. When conditioned on ${\cal G}_{t}$, the entries of both $\bm{Q}_{t}$ and $\bm{R}_{t}$ become deterministic values, leading to the following lemma.
@@
The proof of Lemma 9.10 relies on the rotational invariance of the Gaussian distribution and utilizes Lemma 11 from Bayati and Montanari, (2011). The proofs of this lemma and the subsequent lemma— which describes the distribution of $\vec{\Upsilon}_{t+1}{}$ given the event ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$— follow a similar approach to the proofs of Lemmas 2 and 3 in Shirani and Bayati, (2024), and we refer readers to that work for detailed derivations.
@@
Here, we first state Lemma 9.12 which is an expanded version of Theorem 9.8. To this end, we need some new notations. Specifically, for vectors $\vec{u},\vec{v}\in\mathbb{R}^{m}$, we define the scalar product $\vec{u},\vec{v}\right>:=\frac{1}{m}\sum_{i=1}^{m}u_{i}v_{i}$. Also, considering (27), for $t\geq 1$, we define
@@
For a fixed $k\geq 2$ and a specified subpopulation of experimental units $\mathcal{S}$, consider the following conditions. Given that Assumption 9.4 holds, and both the subpopulation $\mathcal{S}$ and the complete experimental population satisfy Assumption 9.4, the following statements are valid for all time steps $t$ under the BSE equations defined in (27):
@@
    =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big{[}\psi\big{(}Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\tilde{\rho}_{t+1}{}Z_{t}+H^{\mathcal{S}}_{t}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big{)}\Big{]},$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{r+1}{})^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma^{2}_{r}}{N}\sum_{i=1}^{N}(U^{i}_{r}{})^{2}$ (36a)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\Upsilon^{i}_{s+1}{}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (36b)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}^{2}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2},$ (36c)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big{(}\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big{)}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (36d)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\big{(}\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big{)}$ =$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}.$ (36e)
@@
In the following section, we will provide a comprehensive explanation of the conditioning technique, which will be employed to establish the results presented in Lemma 9.12.

### v2

@@ 10.4 Batch-level State Evolution
    Let ${M}^{i}\sim p_{\mu^{i}}$ and ${M}_{t}^{i}\sim p_{\mu^{i}_{t}}$; then, ${M}^{i}$ and ${M}_{t}^{i}$ are independent of all other randomnesses in the model. Additionally, denoting $\bar{\mu}^{i}=\mathbb{E}[{M}^{i}]$ and $\bar{\mu}^{i}_{t}=\mathbb{E}[{M}_{t}^{i}]$, let $p_{\mu}^{\mathcal{S}}(N)$ and $p_{\mu_{t}}^{\mathcal{S}}(N)$ be respectively the empirical distributions of $\bar{\mu}^{i}$ and $\bar{\mu}^{i}_{t}$ with $i\in\mathcal{S}$ and $i\leq N$. Then, $p_{\mu}^{\mathcal{S}}(N)$ and $p_{\mu_{t}}^{\mathcal{S}}(N)$ converge weakly to probability measures $p_{\mu}^{\mathcal{S}}$ and $p_{\mu_{t}}^{\mathcal{S}}$, respectively.
  6. (vi)
    For all $i$ and $t\in[T]_{0}$, suppose that
    $\left(Y^{\mathcal{S}}_{0}{}(N),{\vec{X}}^{\mathcal{S}}(N),\vec{W}^{\mathcal{S}}(N),{M}^{i}(N),{M}_{t}^{i}(N),{M}^{\mathcal{S}}(N),{M}_{t}^{\mathcal{S}}(N)\right)^{\top}$
    $\sim p_{y}^{\mathcal{S}}(N)\times p_{x}^{\mathcal{S}}(N)\times p_{w}^{\mathcal{S}}(N)\times p_{\mu^{i}}(N)\times p_{\mu_{t}^{i}}(N)\times p_{\mu}^{\mathcal{S}}(N)\times p_{\mu_{t}}^{\mathcal{S}}(N),$
    as well as
    $(Y^{\mathcal{S}}_{0}{},{\vec{X}}^{\mathcal{S}},\vec{W}^{\mathcal{S}},{M}^{i},{M}_{t}^{i},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}})^{\top}\sim p_{y}^{\mathcal{S}}\times p_{x}^{\mathcal{S}}\times p_{w}^{\mathcal{S}}\times p_{\mu^{i}}\times p_{\mu_{t}^{i}}\times p_{\mu}^{\mathcal{S}}\times p_{\mu_{t}}^{\mathcal{S}}$
    Then, as $N\rightarrow\infty$, we have
    $\mathbb{E}\left[{\left\|\left(Y^{\mathcal{S}}_{0}{}(N),{\vec{X}}^{\mathcal{S}}(N),\vec{W}^{\mathcal{S}}(N),{M}^{i}(N),{M}_{t}^{i}(N),{M}^{\mathcal{S}}(N),{M}_{t}^{\mathcal{S}}(N)\right)^{\top}\right\|}^{k}\right]$
    $\rightarrow$~ $\mathbb{E}\left[{\left\|(Y^{\mathcal{S}}_{0}{},{\vec{X}}^{\mathcal{S}},\vec{W}^{\mathcal{S}},{M}^{i},{M}_{t}^{i},{M}^{\mathcal{S}},{M}_{t}^{\mathcal{S}})^{\top}\right\|}^{k}\right]<\infty.$
Assumption 10.4 is a standard assumption in statistical theory, ensuring that the empirical distributions of system attributes remain stable and do not diverge as the sample size increases. This assumption holds, for example, when units’ attributes $\left\{(Y^{i}_{0}{},{\vec{X}}^{i},\vec{W}^{i},\vec{\mu}^{\;i\cdot},\vec{\mu}^{\;i\cdot}_{t})\right\}_{i}$ follow an i.i.d. distribution with finite moments of order $k$. Moreover, a wide range of treatment assignments satisfies the conditions of Assumption 10.4, including cases where the support of $\Pi$ (as specified by Remark 3.1) is bounded, such as the Bernoulli design. By imposing such conditions, we ensure the reliability of estimation by keeping the experimental design moments finite and manageable.
To be more specific, Parts (iv) and (v) of Assumption 10.4 formalize and extend Assumption 3.3 from the main text. Indeed, we can simplify the framework by dropping Part (v) under the assumption that $p_{\mu^{i}}$ and $p_{\mu_{t}^{i}}$ are identical across all units $i$. Figure 28 illustrates these assumptions by showing how the underlying $N\times N$ interference matrices must satisfy certain convergent weak limits across both their columns and rows as the population size $N$ increases.
[…]
Figure 28: Conceptual illustration of interference matrix convergence assumptions. The diagram shows how elements of the $N\times N$ interference matrices must exhibit consistent distributional behavior across rows (units) and columns (interactions) as $N\to\infty$, ensuring that both individual unit interference patterns and cross-unit interaction structures converge to well-defined limiting distributions.
Example 10.4
Consider the fixed interference matrix (the left matrix in Figure 28) in a simple setting with two interference levels: $\mu^{ij}\in\{0.3,0.8\}$. For the sequence of settings illustrated in Figure 29, the random variable M{M} follows the distribution $\mathbb{P}({M}=0.3)=\frac{2}{3}$ and $\mathbb{P}({M}=0.8)=\frac{1}{3}$, with ${M}^{i}={M}$ for all units $i$.
[…]
Figure 29: A sequence of interference matrices exhibiting the same interference distribution pattern.
@@
We can drop Assumptions 8 and 10.4-(vi) by confining the functions $g_{t}$ and $h_{t}$ to be bounded and continuous.
Remark 10.6 (Time-dependent covariates)
For each time period $t=0,1,\ldots,T$, let ${\cal X}_{t}\in\mathbb{R}^{n^{\text{x}}_{t}\times N}$ denote the time-dependent covariate matrix, where the $i^{th}$ column corresponds to the covariates for unit $i$ at time $t$. We extend the original covariate matrix $\bm{X}$ by incorporating these time-dependent covariates, resulting in an augmented covariate matrix:
$\bm{X}^{\top}|{\cal X}_{0}^{\top}|\ldots|{\cal X}_{T-1}^{\top}]^{\top}\in\mathbb{R}^{(n^{\text{x}}+n^{\text{x}}_{0}+\ldots+n^{\text{x}}_{T})\times N}.$
The functions $g_{t}$ and $h_{t}$ are then modified to reference the appropriate portion of this extended covariate matrix at each time period. As an example and to streamline subsequent analysis, we exclude the noise vectors $\vec{\epsilon}^{\;}_{t}$ from discussions of the potential outcome specification in Eq. (20) by constructing a new covariate matrix that absorbs the noise terms:
$\bm{X}^{\text{new}}:=[\bm{X}^{\top}|\vec{\epsilon}^{\;}_{0}|\ldots|\vec{\epsilon}^{\;}_{T-1}]^{\top}\in\mathbb{R}^{(n^{\text{x}}_{t}+T)\times N}.$
We then redefine the outcome functions as $g_{t}^{\text{new}}(\cdot,\cdot,\bm{X}^{\text{new}}):=g_{t}(\cdot,\cdot,\bm{X})$ and $h_{t}^{\text{new}}(\cdot,\cdot,\bm{X}^{\text{new}}):=h_{t}(\cdot,\cdot,\bm{X})+\vec{\epsilon}^{\;}_{t}$.
@@
Whenever $\mathcal{S}$ represents the entire experimental population, we omit the superscript $\mathcal{S}$ from all notations.
@@
$\tilde{\nu}_{1}^{\mathcal{S}}{}$ $=\mathbb{E}\left[({M}^{\mathcal{S}}+{M}_{0}^{\mathcal{S}})g_{0}\big(Y_{0}{},\vec{W},{\vec{X}}\big)\right],\quad(\tilde{\rho}_{1}{})^{2}=(\sigma^{2}+\sigma_{0}^{2})\mathbb{E}\left[g_{0}\big(Y_{0}{},\vec{W},{\vec{X}}\big)^{2}\right],\quad H^{\mathcal{S}}_{0}{}=h_{0}\big(Y^{\mathcal{S}}_{0}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big),$ (32)
@@
  • ${M}^{\mathcal{S}}\sim p_{\mu}^{\mathcal{S}}$ and ${M}_{t}^{\mathcal{S}}\sim p_{\mu_{t}}^{\mathcal{S}}$ represent the weak limits of interference elements at the subpopulation level, as specified in Assumption 10.4;
  • $Z_{t}$ follows a standard Gaussian distribution;
  • and $\tilde{\nu}_{t}{}$ and $H_{t}{}$ represent the same quantities as $\tilde{\nu}_{t}^{\mathcal{S}}{}$ and $H^{\mathcal{S}}_{t}{}$, respectively, but for $\mathcal{S}$ to include the entire population (see Remark 10.7).
Remark 10.8
While we focus on the special case of the general model in (20) and (21) with only one lag term $\vec{Y}_{t}{}$, we can readily extend (32) to incorporate more lag terms. For instance, to include an additional lag term in function $g_{t}$, we would modify (32) by including $\tilde{\nu}_{t-1}{}+\tilde{\rho}_{t-1}{}Z_{t-1}+H_{t-2}{}$ as an argument in $g_{t}$. We also need to revise the initial conditions (first line in (32)) according to the number of lag terms.
Now, the following theorem characterizes the distribution of units’ outcomes $Y^{i}_{1}{},\ldots,Y^{i}_{t+1}{}$ within the large sample asymptotic, based on the BSE equations outlined in Eq. (32).
@@
Fixing $k\geq 2$, consider the sequence of units’ attributes $\left\{(Y^{i}_{0}{},{\vec{X}}^{i},\vec{W}^{i},\vec{\mu}^{\;i\cdot},\vec{\mu}^{\;i\cdot}_{t})\right\}_{i,t}$ and suppose that Assumptions 10.2 and 8 hold, and Assumption 10.4 is satisfied once for $\mathcal{S}$ and once for the enitre population. Then, in view of the BSE equations given in (32), for any function $\psi\in{{\cal C}{\cal P}}(k)$, we have,
@@
$\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}$ $\mathbb{E}\Big[\psi\big(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{T}^{\mathcal{S}}{}+\tilde{\rho}_{T}{}Z_{T}+H^{\mathcal{S}}_{T-1}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big)\Big],$
where $Z_{t}\sim{\cal N}(0,1),\;t=1,\ldots,T,$ is independent of $\vec{W}^{\mathcal{S}}\sim p_{w}^{\mathcal{S}}$ and ${\vec{X}}^{\mathcal{S}}\sim p_{x}^{\mathcal{S}}$.
@@
To derive the result of Theorem 4.2 from Theorem 10.9, we note that when $\mathcal{S}$ depends solely on the treatment allocation, which is independently distributed from all other variables, $\mathcal{S}$ effectively functions as a random sample from the experimental population. As a result, the distributions of $Y^{\mathcal{S}}_{0}{}$, ${\vec{X}}^{\mathcal{S}}$, ${M}^{\mathcal{S}}$, and ${M}_{t}^{\mathcal{S}}$ are equivalent to those of $Y_{0}{}$, $\vec{X}$}, M{M}, and ${M}_{t}$, respectively. This equivalence follows from the fact that the empirical distribution of quantities in a random sample converges to the empirical distribution of the entire population.
@@ 10.5 Conditioning Technique
$\bm{Q}_{t}=\left[\vec{U}_{0}{}\Big|\vec{U}_{1}{}\Big|\ldots\Big|\vec{U}_{t-1}{}\right],\quad\quad\bm{R}_{t}=\left[\vec{\Upsilon}_{1}{}-{\bf B}_{0}\vec{U}_{0}{}\Big|\ldots\Big|\vec{\Upsilon}_{t}{}-{\bf B}_{t-1}\vec{U}_{t-1}{}\right].$ (34)
According to Eq. (34), $\bm{Q}_{t}$ and $\bm{R}_{t}$ are matrices with columns of $\vec{U}_{s-1}{}$ and $\vec{\Upsilon}_{s}{}-{\bf B}_{s-1}\vec{U}_{s-1}{}$, when $s=1,\ldots,t$, respectively. Then, we denote by $\vec{U}^{\parallel}_{t}{}$ the projection of $\vec{U}_{t}{}$ onto the space generated by the columns of $\bm{Q}_{t}$ and define $\vec{U}^{\perp}_{t}{}=\vec{U}_{t}{}-\vec{U}^{\parallel}_{t}{}$. Assuming $\bm{Q}_{t}$ is a full-row rank matrix (we prove this is indeed the case in Lemma 10.13-(c)), we also define $\vec{\gamma}_{t}=(\gamma^{0}_{t},\gamma^{1}_{t},\ldots,\gamma^{t-1}_{t})^{\top}$ such that
@@
Now, note that the available observation at any time $t$ implicitly reveals information about the fixed interference matrix $\bf A$}. To manage this intricate randomness, we define ${\cal G}_{t}$ as the $\sigma$-algebra generated by $\vec{Y}_{0}{}$, $\vec{\Upsilon}_{1}{},\ldots,\vec{\Upsilon}_{t}{}$, ${\bf M}_{t}$, ${\bf B}_{0},\ldots,{\bf B}_{t-1}$, $\bm{W}{}$, and $\bm{X}$. We then compute the conditional distribution of $\bf A$} given ${\cal G}_{t}$. Conditioning on ${\cal G}_{t}$, by definition and based on Eq. (21), is equivalent to conditioning on the following event:
$\{{\bf A}\vec{U}_{0}{}=\vec{\Upsilon}_{1}{}-{\bf B}_{0}\vec{U}_{0}{},\;{\bf A}\vec{U}_{1}{}=\vec{\Upsilon}_{2}{}-{\bf B}_{1}\vec{U}_{1}{},\;\ldots,\;{\bf A}\vec{U}_{t-1}{}=\vec{\Upsilon}_{t}{}-{\bf B}_{t-1}\vec{U}_{t-1}{}\}.$
which encapsulates all information we have about the fixed interference matrix $\bf A$} up to time $t$. By Eq. (34), this is also equivalent to the event ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$. When conditioned on ${\cal G}_{t}$, the entries of both $\bm{Q}_{t}$ and $\bm{R}_{t}$ become deterministic values, leading to the following lemma.
@@
The proof of Lemma 10.11 relies on the rotational invariance of the Gaussian distribution and utilizes Lemma 11 from Bayati and Montanari, (2011). The proofs of this lemma and the subsequent lemma, which describes the distribution of $\vec{\Upsilon}_{t+1}{}$ given the event ${\bf A}\bm{Q}_{t}=\bm{R}_{t}$, follow a similar approach to the proofs of Lemmas 2 and 3 in Shirani and Bayati, (2024), and we refer readers to that work for detailed derivations.
@@
We begin by stating Lemma 10.13, which provides an expanded version of Theorem 10.9. This expanded formulation is necessary due to our inductive proof technique, which requires establishing that $\bm{Q}_{t}$ is a full-row rank matrix alongside several auxiliary results needed to advance each step of the induction. To this end, we need some new notations. Specifically, for vectors $\vec{u},\vec{v}\in\mathbb{R}^{m}$, we define the scalar product $\left<\vec{u},\vec{v}\right>:=\frac{1}{m}\sum_{i=1}^{m}u_{i}v_{i}$. Also, considering (32), for $t\geq 1$, we define
@@
For a fixed $k\geq 2$ and a specified subpopulation of experimental units $\mathcal{S}$, suppose that the conditions of Theorem 10.9 hold. Then, the following statements are valid for all time steps $t$ under the BSE equations defined in (32):
@@
    $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\mathbb{E}\Big[\psi\big(Y^{\mathcal{S}}_{0}{},\tilde{\nu}_{1}^{\mathcal{S}}{}+\tilde{\rho}_{1}{}Z_{1}+H^{\mathcal{S}}_{0}{},\ldots,\tilde{\nu}_{t+1}^{\mathcal{S}}{}+\tilde{\rho}_{t+1}{}Z_{t}+H^{\mathcal{S}}_{t}{},\vec{W}^{\mathcal{S}},{\vec{X}}^{\mathcal{S}}\big)\Big],$
@@
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}(\Upsilon^{i}_{r+1}{})^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\tilde{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}+\sigma^{2}_{r}}{N}\sum_{i=1}^{N}(U^{i}_{r}{})^{2}$ (41a)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\Upsilon^{i}_{s+1}{}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (41b)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big)^{2}$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;(\bar{\rho}_{r+1}{})^{2}\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}\left(U^{i}_{r}{}\right)^{2},$ (41c)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\big(\Upsilon^{i}_{r+1}{}-\vec{\text{B}}^{i\cdot}_{r}\vec{U}_{r}{}\big)\big(\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big)$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{},$ (41d)
    $\lim_{N\rightarrow\infty}\frac{1}{N}\sum_{i=1}^{N}\Upsilon^{i}_{r+1}{}\big(\Upsilon^{i}_{s+1}{}-\vec{\text{B}}^{i\cdot}_{s}\vec{U}_{s}{}\big)$ $\;\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize a.s.}}}}{{=}}\;\lim_{N\rightarrow\infty}\frac{\sigma^{2}}{N}\sum_{i=1}^{N}U^{i}_{r}{}U^{i}_{s}{}.$ (41e)